OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2 x + sin x = 0 pleaseee

OpenStudy (anonymous):

sin2x=-sinx I believe this only happens when sin(x)'s are equal to zero. So your answers would be 0, pi

OpenStudy (anonymous):

thats not one of the answers.. theyre 0,π, π/3, 5π/3 0,π,π/3,2π/3 0,π,4π/3,5π/3 0,π,3π/2

OpenStudy (anonymous):

Since you have a multiple choice problem, I would just plug the possible values they give you into the equation and see if they work.

OpenStudy (anonymous):

Factor out a sin(x) giving: \[\sin(x)(\sin(x)+1)=0 \implies \sin(x)=0 or \sin(x)=-1 \implies\]\[ x=n \pi, x= \frac{3\pi}{2},\frac{7 \pi}{2}, \frac{11 \pi}{2}\] So on \[[ 0, 2 \pi) \implies x=0,x=\pi,x=\frac{3 \pi}{2}\]

OpenStudy (anonymous):

omg thanks so much